### v2

@@ Abstract
In recent years, the holographic duality between $T\bar{T}$-deformed conformal field theory (CFT) and Anti-de Sitter (AdS) space with finite radial truncation has received significant attention. The study of $T\bar{T}$ deformation within the framework of de Sitter (dS)/CFT duality has also progressed. This paper shows that the trace flow equation in dS spacetime can be analytically extended from its AdS counterpart through double Wick rotations. Additionally, we generalize the replica sphere method in both AdS and dS holography to derive a general expression for the entanglement entropy of arbitrary single spatial intervals. For both finite size and finite temperature systems, we obtain the entanglement entropy corrections due to $T\bar{T}$ deformation. Finally, using strong subadditivity and boosted strong subadditivity, we show that the dual field theory on the timelike boundary is non-local in dS holography.
@@ 1 Introduction
From the perspective of the AdS/CFT duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , the $T\bar{T}$-deformed CFT corresponds to the semi-classical gravity theory living in an AdS spacetime with finite cutoff McGough:2016lol ; Kraus:2018xrn , as shown in figure 1.
@@
In refs. McGough:2016lol ; Kraus:2018xrn , the authors calculated the propagation speed of the massless modes, energy spectrum, and partition functions in both the contexts of the $T\bar{T}$-deformed quantum field theory and AdS gravity with a finite cutoff. The generic holographic correlation functions have been established in the refs. He:2023hoj ; He:2023knl ; He:2024fdm ; He:2024xbi . These results support the duality depicted in figure 1. In addition to the aforementioned correspondences, another significant result of holographic principles is that the entanglement entropy of the boundary field theory equals the area of the minimal surface, also known as the Ryu-Takayanagi (RT) surface, inside the AdS spacetime Hubeny:2007xt ; Ryu:2006bv . The RT formula for holographic entanglement entropy serves as a criterion for validating the holographic principle, with $T\bar{T}$-deformed theories offering an additional test. For the holographic duality with $T\bar{T}$ deformation, this duality has been discussed in refs. Chen:2018eqk ; Jeong:2019ylz ; Jiang:2019epa ; Allameh:2021moy ; Setare:2022qls ; He:2023xnb ; Apolo:2023ckr . In ref. Chen:2018eqk , the authors investigated the entanglement entropy of the $T\bar{T}$-deformed field theory at the level of perturbation (2). They found that for the system at finite temperature, the entanglement entropy receives a first-order $\lambda$ correction from the $T\bar{T}$ deformation, but for a system of finite size, this first-order correction disappears. To explore the impact of $T\bar{T}$ deformation on the entanglement entropy in finite-size systems, ref. Jiang:2023ffu examined the timelike entanglement entropy, drawing similar conclusions to previous studies Li:2022tsv ; Doi:2023zaf . They found that the timelike entanglement entropy receives no first-order correction at finite temperature, but there is a first-order correction in the finite size case. Additionally, the reflected entropy Dutta:2019gen and pseudo entropy Nakata:2020luh of the $T\bar{T}$ deformed field theory have been studied in refs. Basu:2024bal ; He:2023wko . For more interesting work on $T\bar{T}$ deformation, please refer to refs. He:2019vzf ; Asrat:2020uib ; Khoeini-Moghaddam:2020ymm ; Li:2020pwa ; Jeong:2022jmp ; He:2023obo ; Tian:2023fgf ; Apolo:2023vnm ; Basu:2023bov ; Basu:2023aqz ; He:2024pbp ; Tsolakidis:2024wut ; Babaei-Aghbolagh:2024hti .
The holographic principle has been well established in the AdS spacetime with finite cutoff $r=r_{c}$. However, given that our real universe is more accurately described by a dS metric, it is crucial to explore whether the holographic principle extends to the dS spacetime Witten:2001kn ; Maldacena:2002vr ; Miyaji:2015yva ; Jiang:2023loq ; Chang:2023gkt ; Chen:2023prz ; Doi:2024nty . In the usual dS/CFT duality Strominger:2001pn ; Strominger:2001gp , the CFT resides on a future spacelike boundary, which renders the dual field theory non-unitary Maldacena:2002vr ; Doi:2022iyj . Due to the challenges of the dS/CFT duality Anninos:2011ui ; Ng:2012xp , there are only a few clearly established examples Hikida:2022ltr ; Hikida:2021ese ; Chen:2022ozy . In recent years, other holographic duality models in the dS spacetime, such as the dS/dS correspondence Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Geng:2019ruz ; Geng:2020kxh ; Geng:2021wcq and half-dS holography Kawamoto:2023nki , have gained attention and been studied. In these two models, the holographic boundary is a timelike boundary of the dS spacetime, so the dual field theory is expected to be unitary. In the dS/dS correspondence, the cosmological constant term also appears in the trace flow equation of the $T\bar{T}$ deformation, thus the $T\bar{T}$ deformation of this model is referred to as $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Batra:2024kjl .
@@
We aim to study the impact of $T\bar{T}$ deformation on the entanglement entropy in the dS/CFT duality. However, the deformed field theory no longer retains conformal symmetry, rendering the original method for calculating the entanglement entropy inapplicable. Fortunately, when the deformed CFT is situated on a two-dimensional sphere, the entanglement entropy of a subregion with endpoints at antipodal points can be calculated via the replica method Calabrese:2009qy ; Donnelly:2018bef ; Caputa:2019pam ; Deng:2023pjs . In this context, by adjusting the spherical radius $r$, we can study the renormalization group flow Wilson:1973jj ; Polchinski:1983gv of the dual field theory under $T\bar{T}$ deformation, with $r$ functioning as the radial coordinate in the bulk spacetime. This method is an extension of the approach proposed in ref. Lewkowycz:2013nqa and was further generalized in ref. Lewkowycz:2019xse to compute the entanglement entropy of arbitrary spatial intervals on a two-dimensional plane. In ref. Chen:2023eic , the replica method was employed within the $T\bar{T}$-deformed dS/CFT duality to calculate the entanglement entropy for the subregion between two antipodal points on the sphere.
@@
In this paper, we present a general expression for the entanglement entropy of arbitrary spatial intervals in both AdS and dS holography by combining the Casini-Huerta-Myers (CHM) map Casini:2011kv with the replica method on a sphere. The results are
@@
  • AdS($0<r_{\text{eff}}<\infty$): $S=\frac{c_{\rm AdS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\text{eff}}^{2}}}}\right),$
@@
  • dS($\ell_{\rm dS}<r_{\text{eff}}<\infty$): $S=-\frac{ic_{\rm dS}}{3}\mathrm{arccoth}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$
  • dS($0<r_{\text{eff}}<\ell_{\rm dS}$): $S=\frac{c_{\rm dS}}{3}\arctan\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right),$
@@
$S=-\frac{ic_{\rm dS}}{3}\mathrm{arctanh}\,\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)\,.$ (3)
When the effective radius of the sphere is greater than the cosmological horizon radius ($\ell_{\rm dS}<r_{\text{eff}}<\infty$), this formula corresponds to the result of the $T\bar{T}$-deformed dS/CFT duality, where the dual field theory is non-unitary. When the effective radius of the sphere is less than the cosmological horizon radius ($0<r_{\text{eff}}<\ell_{\rm dS}$), this formula agrees with the results of dS/dS correspondence and half-dS holography, where the dual field theory is non-local Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki .
This paper is organized as follows: In Section 2, we review the results from ref. Kraus:2018xrn to derive the trace flow equations in AdS/CFT with Euclidean signature and in dS/CFT with Lorentzian signature. We then demonstrate the extension from one result to the other via a double Wick rotation. In Section 3, we review the calculation of the entanglement entropy of the $T\bar{T}$-deformed CFT, providing general expressions for both AdS and dS holographies. Based on these findings, we calculate the $C$-function and discuss the non-unitarity and non-locality of the dual field theory. In Section 4, we review and generalize the CHM method for calculating the entanglement entropy of arbitrary spatial intervals in the $T\bar{T}$-deformed theory. We explore the effects of the $T\bar{T}$ deformation in finite temperature and finite size systems. In Section 5, we first verify the calculation of the entanglement entropy in the AdS/dS correspondence. We then examine the non-locality of the dual field theory in dS/dS correspondence and half-dS holography by analyzing the strong subadditivity and the boosted strong subadditivity. Finally, we summarize our results and present a discussion in Section 6.
@@ 2 Holographic $T\bar{T}$ Flow
In this section, we review the holographic derivation of the trace flow equation for the $T\bar{T}$ deformation Shyam:2017znq ; Hartman:2018tkw . We then show that the trace flow equation in the dS/CFT duality can be obtained from its AdS counterpart via a double Wick rotation.
@@ 2.1 $T\bar{T}$ Flow in AdS/CFT
The EAdS3 metric in Fefferman-Graham Fefferman1985 coordinates is
@@ 2.2 $T\bar{T}$ Flow in dS/CFT
where $T^{ij}=-\frac{2i}{\sqrt{\gamma}}\Pi^{ij}$. This leads to the $T\bar{T}$ trace flow equation for the dS/CFT is
@@
By comparing this with the field theory equation, the correspondences between the Brown-Henneaux central charge, the deformation parameter, and the gravitational quantities for the dS spacetime Chen:2023eic are given by
$c=-i\frac{3\ell_{\rm AdS}}{2G}=-ic_{\rm dS}\,,\quad\lambda=-i8G\ell_{\rm AdS}=-i\lambda_{\rm dS}\,.$ (23)
@@ 2.3 Double Wick Rotation to dS/CFT
$S_{\rm AdS}$ $\rightarrow-\frac{i}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)$
$\quad-\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(iK-\frac{i}{\ell_{\rm dS}}\right)$
@@
The double Wick rotation implies that the partition function of AdS gravity corresponds to the Bunch-Davies wavefunction of dS gravity. Additionally, all quantities in the Wheeler-DeWitt equation for the AdS spacetime are mapped to their corresponding quantities in the Wheeler-DeWitt equation for the dS spacetime. As a result, eqs. (12) and (14) correspond to eqs. (20) and (21), respectively. Therefore, the trace flow equation in the dS/CFT can be directly derived by applying this rotation to the trace flow equation in the AdS/CFT.
@@ 3 The Replica Method for the $T\bar{T}$-Deformed Field Theory
The entanglement entropy of the subregion $A$ is computed by taking the limit $n\rightarrow 1$ of the conical entropy Headrick:2010zt ; Hung:2011nu ; Lewkowycz:2013nqa
@@
$\begin{cases}{T^{a}}_{a}=-\frac{c}{24\pi}R^{(2)}-\frac{\pi\lambda}{2}\left[T^{ab}T_{ab}-({T^{a}}_{a})^{2}\right],\\ \nabla_{a}{T^{a}}_{b}=0.\end{cases}$ (31)
@@
In this case, the parameter $\alpha$ determines the stress tensor, which is uniquely defined by the trace flow equations. This approach is specifically designed to calculate the entanglement entropy of the subregion between two antipodal points on the sphere. Future work aims to extend this framework to explore broader implications of the holographic principle, building on the refs. Grieninger:2019zts ; Lewkowycz:2019xse .
3.1 AdS case

### v3

@@ Abstract
In recent years, the holographic duality between $T\bar{T}$-deformed conformal field theory (CFT) and Anti-de Sitter (AdS) spacetime with finite radial cutoff has received significant attention. The study of $T\bar{T}$ deformation within the framework of de Sitter (dS)/CFT duality has also progressed. This paper shows that the trace flow equation in dS spacetime can be analytically extended from its AdS counterpart through double Wick rotations. Meanwhile, we generalize the replica method in both AdS and dS holography to derive a general expression for the entanglement entropy of arbitrary single spatial intervals within the $T\bar{T}$-deformed framework. For both finite size and finite temperature systems, we obtain the analytical expression for the entanglement entropy after $T\bar{T}$ deformation. Finally, in dS/dS holography and half-dS holography, we find that the dual field theory exhibits non-locality by analyzing the strong subadditivity and boosted strong subadditivity of entanglement entropy.
@@ 1 Introduction
From the perspective of the AdS/CFT duality Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj , the $T\bar{T}$-deformed CFT corresponds to the semiclassical gravity theory living in an AdS spacetime with finite cutoff McGough:2016lol ; Kraus:2018xrn , as shown in figure 1.
@@
In refs. McGough:2016lol ; Kraus:2018xrn , the authors calculated the propagation speed of the massless modes, energy spectrum, and partition functions in both the contexts of the $T\bar{T}$-deformed quantum field theory and AdS gravity with a finite cutoff. The generic holographic correlation functions have been established in the refs. He:2023hoj ; He:2023knl ; He:2024fdm ; He:2024xbi . These results support the duality depicted in figure 1. In addition to the aforementioned correspondences, another significant result of holographic principles is that the entanglement entropy of the boundary field theory equals the area of the minimal surface, also known as the Ryu-Takayanagi (RT) surface, inside the AdS spacetime Hubeny:2007xt ; Ryu:2006bv . The RT formula for holographic entanglement entropy serves as a criterion for validating the holographic principle, with $T\bar{T}$-deformed theories offering an additional test. For the holographic duality with $T\bar{T}$ deformation, this duality has been discussed in refs. Chen:2018eqk ; Jeong:2019ylz ; Jiang:2019epa ; Allameh:2021moy ; Setare:2022qls ; He:2023xnb ; Apolo:2023ckr ; Apolo:2023vnm . Meanwhile, the reflected entropy Dutta:2019gen and pseudo entropy Nakata:2020luh of the $T\bar{T}$-deformed field theory have been studied in refs. Basu:2024bal ; He:2023wko . For more interesting work on $T\bar{T}$ deformation, please refer to refs. He:2019vzf ; Asrat:2020uib ; Khoeini-Moghaddam:2020ymm ; Li:2020pwa ; Jeong:2022jmp ; He:2023obo ; Tian:2023fgf ; Basu:2023bov ; Basu:2023aqz ; He:2024pbp ; Tsolakidis:2024wut ; Babaei-Aghbolagh:2024hti ; He:2025ppz ; He:2025fdz .
Although the holographic principle has been well established in the AdS spacetime with finite cutoff $r=r_{c}$, it remains crucial to investigate whether the holographic principle extends to dS spacetime Witten:2001kn ; Maldacena:2002vr ; Anninos:2012ft ; Miyaji:2015yva ; Narayan:2017xca ; Hikida:2021ese ; Chen:2022ozy ; Chang:2023gkt ; Chen:2023prz ; Doi:2024nty , since our universe is more accurately described by a dS metric. In the standard dS/CFT duality Strominger:2001pn ; Strominger:2001gp , the CFT resides at a future spacelike infinite, which leads to some exotic characteristics Anninos:2011ui ; Ng:2012xp ; Hikida:2022ltr ; Doi:2022iyj . For example, the central charge is imaginary. Based on these exotic characteristics, a precise construction of the dual CFT in dS spacetime remains elusive. For the holographic entanglement entropy in dS/CFT, by calculating the extreme value of the area of the RT surface anchored at the future spacelike infinite boundary on the gravitational side, refs. Narayan:2015vda ; Narayan:2015oka ; Narayan:2016xwq ; Narayan:2022afv ; Narayan:2023zen have revealed the presence of complex-valued saddle points, which coincides with the imaginary central charge. At the same time, this observation aligns with results obtained from the analytic continuation of holographic entanglement entropy in Euclidean AdS spacetime, thereby offering a potential avenue to explore the dS/CFT duality through double Wick rotation Sato:2015tta . In addition, other holographic duality models in the dS spacetime, such as the dS/dS holography Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Geng:2019ruz ; Geng:2020kxh ; Geng:2021wcq and half-dS holography Kawamoto:2023nki , have also gained attention and been studied. In these two models, the dual field theory resides on the timelike boundary of the dS spacetime.
@@
The primary purpose of this paper is to study the impact of the $T\bar{T}$ deformation on holographic entanglement entropy in the dS spacetime, including three holographic models: dS/CFT, dS/dS, and half-dS. For the $T\bar{T}$-deformed dS/CFT Chen:2023eic , the dual field theory is on the spacelike slice $t=t_{c}$, which coincides with the idea of Cauchy slice holography Araujo-Regado:2022gvw ; Araujo-Regado:2022jpj . In the $T\bar{T}$-deformed dS/dS, the cosmological constant term also appears in the trace flow equation of the $T\bar{T}$ deformation, thus the $T\bar{T}$ deformation of this model is referred to as $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Batra:2024kjl . However, the deformed field theory no longer retains conformal symmetry, rendering the original method for calculating the entanglement entropy inapplicable. Fortunately, when the $T\bar{T}$-deformed CFT is situated on a two-dimensional sphere, the entanglement entropy of a subregion with endpoints at antipodal points can be calculated by replicating the spherical geometry Calabrese:2009qy ; Donnelly:2018bef ; Caputa:2019pam ; Deng:2023pjs ; Grieninger:2019zts . In this context, by adjusting the spherical radius $r$, we can also study the renormalization group flow Wilson:1973jj ; Polchinski:1983gv of the dual field theory under $T\bar{T}$ deformation, with $r$ functioning as the radial coordinate in the bulk spacetime. This method is an extension of the approach proposed in ref. Lewkowycz:2013nqa and was further generalized in ref. Lewkowycz:2019xse to compute the entanglement entropy of arbitrary spatial intervals on a two-dimensional plane. In ref. Chen:2023eic , the replica method was employed within the $T\bar{T}$-deformed dS/CFT duality to calculate the entanglement entropy for the subregion between two antipodal points on the sphere. In the $T\bar{T}$-deformed theories, non-perturbative calculations of entanglement entropy have been achieved using the replica trick, yielding exact results on spherical and flat manifolds Donnelly:2018bef ; Lewkowycz:2019xse . However, the non-perturbative calculation on general manifolds is still lacking.
@@
In this paper, we present a general expression for the entanglement entropy of arbitrary spatial intervals in general manifolds by combining the Casini-Huerta-Myers (CHM) map Casini:2011kv with the replica method on a sphere. The results are
@@
  • AdS($0<r_{\text{eff}}<\infty$): $S=\frac{c_{\rm AdS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1+\frac{c_{\rm AdS}\lambda_{\rm AdS}}{12r_{\text{eff}}^{2}}}}\right),$
@@
  • dS($\ell_{\rm dS}<r_{\text{eff}}<\infty$): $S=-\frac{ic_{\rm dS}}{3}\coth^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)+\frac{\pi c_{\rm dS}}{6},$
  • dS($0<r_{\text{eff}}<\ell_{\rm dS}$): $S=\frac{c_{\rm dS}}{3}\tan^{-1}\left(\frac{1}{\sqrt{\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}-1}}\right),$
@@
$S=-\frac{ic_{\rm dS}}{3}\tanh^{-1}\left(\frac{1}{\sqrt{1-\frac{c_{\rm dS}\lambda_{\rm dS}}{12r_{\text{eff}}^{2}}}}\right)\,.$ (3)
When the effective radius of the sphere is greater than the cosmological horizon radius ($\ell_{\rm dS}<r_{\text{eff}}<\infty$), this formula corresponds to the result of the $T\bar{T}$-deformed dS/CFT duality. When the effective radius of the sphere is less than the cosmological horizon radius ($0<r_{\text{eff}}<\ell_{\rm dS}$), this formula agrees with the results of dS/dS holography and half-dS holography. Through an analysis of entanglement entropy compliance with strong subadditivity (SSA) and boosted strong subadditivity (BSSA) in both dS/dS holography and half-dS holography, we demonstrate that the dual field theory--residing on a timelike slice in dS spacetime--exhibits nonlocality in these holographic frameworks Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki .
This paper is organized as follows: In Section 2, we review the derivation of trace flow equations in AdS/CFT with Euclidean signature and in dS/CFT with Lorentzian signature. Subsequently, we demonstrate the analytic continuation from the AdS/CFT result to the dS/CFT counterpart through double Wick rotation. In Section 3, we review the computation of entanglement entropy for $T\bar{T}$-deformed CFTs, presenting general expressions applicable to both AdS and dS holography. In Section 4, we calculate the non-perturbative results for the entanglement entropy of arbitrary single spatial intervals in $T\bar{T}$-deformed theories on general manifolds by employing the CHM map. We further investigate the impact of $T\bar{T}$ deformations in finite temperature and finite size systems. In Section 5, we examine the non-locality of the dual field theory in dS/dS and half-dS holography by scrutinizing SSA and BSSA. Finally, we summarize our findings and discuss potential future directions in Section 6.
@@ 2 Holographic $T\bar{T}$ Flow
In this section, we review the holographic derivation of the trace flow equation for the $T\bar{T}$ deformation Shyam:2017znq ; Hartman:2018tkw ; Kraus:2018xrn . We then show that the trace flow equation in the dS/CFT duality can be obtained from its AdS counterpart via a double Wick rotation.
@@ 2.1 $T\bar{T}$ Flow in AdS/CFT
The EAdS3 metric in Fefferman-Graham FeffermanGraham1985 coordinates is
@@ 2.2 $T\bar{T}$ Flow in dS/CFT
where $T^{ij}=-\frac{2i}{\sqrt{\gamma}}\Pi^{ij}$. This leads to the $T\bar{T}$ trace flow equation for the dS/CFT being
@@
By comparing this with the field theory equation, the correspondences between the Brown-Henneaux central charge, the deformation parameter, and the gravitational quantities for the dS spacetime are given by
$c=-i\frac{3\ell_{\rm dS}}{2G}=-ic_{\rm dS}\,,\quad\lambda=-i8G\ell_{\rm dS}=-i\lambda_{\rm dS}\,.$ (23)
@@ 2.3 Double Wick Rotation to dS/CFT
$S_{\rm AdS}$ $\rightarrow-\frac{i}{16\pi G}\int_{\mathcal{M}}d^{3}x\sqrt{-g}\left(\mathcal{R}^{(3)}-\frac{2}{\ell_{\rm dS}^{2}}\right)-\frac{1}{8\pi G}\int_{\partial\mathcal{M}}d^{2}x\sqrt{\gamma}\left(iK-\frac{i}{\ell_{\rm dS}}\right)$
@@
Under the saddle point approximation, the double Wick rotation establishes that the wavefunction of AdS gravity, given by $\Psi=e^{-S_{\text{AdS}}}$, corresponds to the Bunch-Davies wavefunction of dS gravity, expressed as $\Psi_{\text{dS}}=e^{iS_{\text{dS}}}$. As a result, eqs. (12) and (14) correspond to eqs. (20) and (21), respectively. This means that the trace flow equation (22) in the dS/CFT can be directly derived by applying the double Wick rotation to the trace flow equation (9) in the AdS/CFT. At the same time, it should be noted that although we have linked the WDW constraints in dS spacetime with the WDW constraints in AdS spacetime through the double Wick rotation, the Hilbert space of dS quantum gravity may not be derived by the analytical continuation from that of AdS [^3] Thanks to the referee for pointing out the limitations of the analytical continuation method here.. This is because the radial wave functional in AdS spacetime obeys regularity at $r=0$, fixing the asymptotic form at $r=\infty$ to the phase factor times the partition function of the CFT. This constraint does not apply in dS spacetimeChakraborty:2023yed ; Cianfrani:2013oja ; Chakraborty:2023los .
@@ 3 The Replica Method for the $T\bar{T}$-Deformed Field Theory
The entanglement entropy of the subregion $A$ is computed by taking the limit $n\rightarrow 1$ of the conical entropy Headrick:2010zt ; Hung:2011nu
@@
$T^{a}_{a}=-\frac{c}{24\pi}R^{(2)}-\frac{\pi\lambda}{2}\left[T^{ab}T_{ab}-(T^{a}_{a})^{2}\right],$ (31)
$\nabla_{a}T^{a}_{b}=0.$ (32)
@@
In this case, the parameter $\alpha$ determines the stress tensor, which is uniquely defined by the trace flow equations.
3.1 AdS spacetime